OpenStudy (anonymous):

In the equation of the circle (x-5)^2+(y-5)^2=50,does it touches with the origin?Find the center of the circle?

OpenStudy (anonymous):

>>In the equation of the circle (x-5)^2+(y-5)^2=50,does it touches with the origin? Put x=0 and y=0 and it's satisfies the equation so it does. >>Find the center of the circle? (5,5)

OpenStudy (campbell_st):

Centre (5,5) so draw a right triangle.... and see if the hypotenuse is equal to the radius.  \[r = 5\sqrt{2}\] from the circle equation using Pythagoras 's theorem.. 5^2 + 5^2 = ch^2 ch ^2 = 50... sam as sth

OpenStudy (campbell_st):

tHE hypotenuse is \[h = 5\sqrt{?}\] the same as the radius so the circle touches the origin

OpenStudy (campbell_st):

? = 2